i. A fixed mass of gas has a pressure of 1.2 MPa and volume of 0.3 m3 and temperature of 100 oC . The pressure is changed to 400 kPa and the volume changed to 0.06 m3.

Determine the new temperature?


1
Expert's answer
2021-05-05T07:41:17-0400
"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2}\\
\frac{1.2\cdot10^6\cdot0.3}{273+100}=\frac{0.4\cdot10^6\cdot0.06}{T_2}\\T_2=-248\degree C"